An ideal air-standard Brayton cycle operates at steady state with compressor inlet conditions of 300K

and 100kPa and a fixed turbine inlet temperature of 1700K. For the cycle, determine (a) the pressure

ratio for given thermal efficiency of 45% and (b) the net power developed


1
Expert's answer
2021-06-16T09:21:38-0400

Answer:-


"T_1=300k \\
p_1=100k\\
T_3=1700k\\"


"r_p=(\frac{P_2}{P_1})=8"

assume air working medium

r=1.4 , "C_p=1.005 K/kg , m=1kg\\"

"process \ \ 1-2\\
\frac{T_2}{T_1}=(\frac{P_2}{P_1})^{1.4-1\over 1.4}"

"T_2=300\times (8)^{1.4-1\over 1.4}"

"\boxed{T_2=543.43K}"


process 3-4

with same formula

"T_4=938.5K"


net power=power in turbine-power in compression

Net Power (P)="mC_p(T_3-T_4)-mC_p(T_2-T_1)"

putting values we get

"P_{net}=1\times1.005[(1700-938.5)]"

=520.66Kw m=1kg/sec


"\eta=1-\frac{1}{(r_p)^{\frac{r-1}{r}}}"

"\eta=1-\frac{1}{(8)^{\frac{1.4-1}{1.4}}}"

"\boxed{\eta_{th}=45 \%}"